Helium gas is contained in a tank with a pressure of 12.4 MPa. If the temperature inside the tank is 32.8 Celsius and the volume of the tank is 18.0 L, determine the mass of the helium in the tank.


1
Expert's answer
2022-03-01T15:37:04-0500

Ideal gas law:

pV = nRT

p = 12.4 MPa = 12.4 x 106 Pa = 122.4 atm

T = 32.8°C = 32.8 + 273.15 = 305.95 K

V = 18.0 L

R = 0.08206 L atm mol−1 K−1

n = pV/RT

n = (122.4 atm)(18.0 L)/(0.08206 L atm mol−1 K−1)(305.95 K) = 87.76 mol He

n = m/Mr

m = n x Mr = 87.76 mol x 4 g/mol = 351.04 g